Think of something that you might want to measure that is affected by random variation.  Identify what you want to measure, then describe its (approximate) sample space. Give a rough description of the probabilities associated with those values (you can simply specify if they are all the same probability or if values in one range will be more likely than values in another range). What would you say to a person who says that he or she "knows" what the outcome of an individual observation will be (an outcome of something that has not happened yet that is subject to random error)?

1
Expert's answer
2021-07-14T11:49:56-0400

For example, if we throw a coin then probabilities of events head, tail will be "\frac{1}{2}, \frac{1}{2}" or if we roll a dice then probability of getting any of 1, 2, 3, 4, 5 and 6 will be "\frac{1}{6}" . Hence, in this example both of cases probability are equal for each event.

But if we have a 4 white ball and 5 black ball in container then probability of drawing white ball is "\frac{4}{9}" and probability of black ball is "\frac{5}{9}" . Therefore in this example probability are not equal for white and black ball.

Hence probability is affected by a random variation.